Skip to Search

 
diagonal spacer

Information Technology Staff Meeting Notes
February 27, 2009

Present: Thomas Bielicke, Cynthia Bowker, Laurie Brennan, Will Chaney, Logan Clapp, Matthew Daley, Debra Fetzer, Jason Fleming, Winston Harris, Paul McDonough, Bobbie Meng, Tom Minton, Chris Nicolich, Amy Polk, Patricia Ruwell, Evan Wack, Stephen Williams
Announcements
  1. Library Council review - budget was the main focus. 23 positions are earmarked for layoff. Everything depends on the legislature. Town Hall meeting today should shed a bit more light. Machen to submit his proposal by April and we should know by June.
  2. Group policy action plan - consolidation of IT across campus - the topic came up at a training session recently. It has been forthcoming for years. AD and Exchange were part of the original component for consolidation. Web and SQL could be next.
  3. HTC300 (the AC in the server room) has a burned out motor. Service call has been placed.
Review Minutes of 02/06/09
  1. Support Issues #7 - Carol was unable to open (not rename) folders.
  2. Hardware #1 - if a CD is dropped in, it will NOT play immediately.
Ongoing IT Updates
Equipment Received/Installed/Repaired
  1. Orders are ongoing.
  2. Loaner laptops - we may be able to replace them. The repair warranty on the student checkout laptops needs to be checked on.
  3. We should have a request from West forthcoming to replace the laptops in West 211 & 212 with desktops.
Support Issues
  1. Recently, there was a ticket request for a second monitor for student use. We are not going to do it. Be sure to clear it with Will first, even if the requester claims Will said to do it.
  2. AD migration issues - student access expiration policy - need to create the student group and include volunteers. Debra to send out notification of account disabling at the end of the spring semester. Summer will be a good time to implement.
  3. Address book reverting - no real fix.
  4. Folder permissions - problems with them as read only. Have to give them full rights on the folder.
  5. ALEPH - can't print. Have to give them permissions.
  6. Admin rights - we aren't allowing anyone to have them. If they need admin rights to a machine, we are thinking of a policy that gives them rights for 20 minutes and then locks them out again.
  7. Logan requests that if asked to pre-stage or rebuild a workstation, please provide the user and location.
  8. Scott Fagen's machine has been retrieved. There are also two machines on the fifth floor to be returned to the IT Dept.
  9. Winston talked to Michelle about possibly moving all of the Ariel stuff to our server. All ILL everywhere will be down for a week, so that would be a good time for the move. Winston and Stephen will meet with ILL.
  10. When a computer is pulled for the vault, be sure the record is removed from AD. A list can be sent to Logan or Amy.
  11. Discussion of inventory techniques used in Grover by the Hardware Support Unit - ways to improve it next year, integrating SMS to distinguish users.
  12. SMS- ongoing. Need to move it to AD. Amy sent an email to the AD folks.
  13. Disabling auto play - still on hold. Will to check with Admin.
  14. EPO - last night, it was only seeing 37 machines instead of 200. Still working on it.
  15. We got a new tape drive for Backup Exec. The heads on the old drive were worn out.
Hardware Support Unit
  1. Steady State has been installed everywhere. Before it reboots in West, it hangs a bit. Problems with one machine in Music and three in MSL - issues with Steady State and Dell. Updated the bios, but that didn't fix it. Deep Freeze wasn't installing the updates. Now they are getting them.
  2. Inventory finished in Music, AFA and West (except Admin). We're about a quarter of the way through with West public. MSL next week, then East.
  3. Inventory issues - conditions in certain offices are hazardous and/or not conducive to doing the job at hand. Let Will know where the conditions exist and he will speak with the appropriate supervisor.
Software Projects Status
  1. Gator Scholar - can be moved to pending outside forces. Val and Sarah are keying data.
  2. ARL Stats ongoing.
  3. Tracking database - Tom is wrapping it up. It can be taken off the list.
  4. Tom is now working on FilmLog. He has been meeting with Laurie Taylor. The first stage is close to done.
  5. UFDC utilization is done.
  6. Affiliation Modification Proposal - the ball is in Bridge's court.
  7. ITS - Stephen and Bonnie - got HSC loaded into the database so they can sign up for classes. Bonnie wants to load historical info, but we can't do it until the HSC integration is complete.
  8. Patron update migration to SQL - Chris has finished it. Should run this weekend and we'll see the results on Monday.
  9. Root certificate for all applications - Logan is working on it.
  10. Blog server reboot brought Grover down briefly this morning. Chris is working to patch Grover so that it doesn't happen again.
  11. NT4 - Chris is working to pull all active databases off of there and moving them to 2K3SQL. NT4 to be re-purposed.
  12. Virtual VM SQL - that's the long range plan. Need to check the licensing. We should probably upgrade to SQL2008.
  13. Success with Shibboleth - Winston and Stephen worked with CNS to get it working. ARPs requested from Bridges.
  14. Servers to move to AD - summer looks like a good time.
  15. ILLIAD billing taskforce - met this week. Stephen has been developing Access reports. Made progress this week. Ongoing.
DLC
  1. Recently rebuilt both Greenstone servers. Problem: Mark can build Lucene index on Development, but not on the Production box. The issue is related to Java. Because the boxes are mirrored now, we need it to work on both servers. Currently, there is no failover backup. May create a natural disk image and reload it on Production.
  2. Tivoli backup is almost in place. Only the file system will be backed up. Once it's complete it will run daily incremental backups.
  3. Hard drives for DLC - received. Fixed the Copibook machine. Should have redundancy.
  4. Server drive is in Logan's office. It should be installed next week.
ALEPH
  1. Carol Whitmer reported a Aleph module crashing. When that happens, you must uninstall & reinstall.
Web Projects
  1. New home page debuted. Upper level pages have been updated. There have not been too many comments. The old page is scheduled to disappear during spring break.
  2. Tom has a rudimentary IPhone page set up, at the request of Admin. Low graphic version.
  3. Usability lab - Tom has requested it to test apps and sites.
Students and Department
  1. Spring break: March 9-13. No students.
IT Security Management
  1. Meeting held.

 

University of Florida Home Page